OpenStudy (anonymous):

31) need answer If 6x - 3y = 30 and 4x = 2 - y then find x + y. -4 -2 0 2 4

OpenStudy (espex):

Set it up as a system of equations, solve for y, solve for x, then add x+y for your answer.

OpenStudy (anonymous):

4x = 2 - y....change to 4x + y = 2 now you system of equations is : 6x - 3y = 30 4x + y = 2 -->(3)4x + y = 2 ------------ 6x - 3y = 30 12x + 3y = 6 (result of multiplying by 3) ------------add 18x + 0 = 36 18x = 36 x = 36/18 x = 2 now sub 2 in for x in either of the original equations 6x - 3y = 30 6(2) - 3y = 30 12 - 3y = 30 -3y = 30 - 12 -3y = 18 y = -18/3 y = -6 Lets check it... 4x + y = 2 4(2) + (-6) = 2 8 - 6 = 2 2 = 2 (correct) now we know that x = 2 and y = - 6. So lets find x + y by subbing in our know values... x + y = 2 + (-6) = 2 - 6 = - 4 --> your answer. Do you have any questions ? If so, please ask and I will be happy to explain.